.\Car A (m = 1500 kg) initially moving at a certain speed to the east collides head on with car B (m = 1000 kg) initially moving to the west at 15 m/s. After the collision, the two cars move together at 2.5 m/s to the east. What is the initial speed of car A?



1
Expert's answer
2021-11-11T17:08:05-0500

"MV-mu=(M+m)v\to 1500\cdot V-1000\cdot15=(1500+1000)\cdot2.5\to"


"V=14.17\ (m/s)" . Answer